OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). ) ?

OpenStudy (anonymous):

i dont understand this question and really need help soo please hurry thanks:)

OpenStudy (anonymous):

ASAP

OpenStudy (anonymous):

PLEASEEEEE

OpenStudy (anonymous):

\[3\sqrt{2x} +x \sqrt{8x} - 5 \sqrt{18x}\] =\[3\sqrt{2x} +2x \sqrt{2x} - 15 \sqrt{2x}\] =\[ 2x \sqrt{2x} - 12 \sqrt{2x}\] =\[ \sqrt{2x} (2x-12)\]

OpenStudy (anonymous):

so technically, everything was wrong